OpenStudy (anonymous):

Sergio’s baby sister gained two pounds each month for her first several months of life. If she weighed 5 pounds when she was born, how many pounds did she weigh after four months? The expression is 5+ 2m , where m is the number of months.

OpenStudy (anonymous):

5+2m = 5+2(4) = 13